OpenStudy (anonymous):

Verify this is an Identity: cos t + cos3t/sin t - sin 3t = -cot t This is what I have so far, cos t + cos3t/sin t - sin 3t = -cot t cos t + (4cos^3 t - 2 cos t)/sin t -(3sin t - 4sin^3 t) = -cot t cos t + cos t(1+3cos^3t - 2 cos t)/sin t - sin(1-2sin t - 3sin^3 t) = -cot t 2cos t (1+3cos^3t - 2 cos t)/ -2sin t(1-2sin t - 3sin^3 t) = -cot t

OpenStudy (anonymous):

I'm confused on what to do next.

OpenStudy (chihiroasleaf):

I think you did some mistakes, \[\cos (3t) = 4 \cos ^{3} (t) - 3 \cos (t)\]

hartnn (hartnn):

alternative way : use SinC - SinD = 2Cos[(C+D)/2]Sin[(C-D)/2] CosC + CosD = 2Cos[(C+D)/2]Cos[(C-D)/2]

hartnn (hartnn):

C=t, D=3t

OpenStudy (chihiroasleaf):

well, I got confused about the problems, which one is the left side?